OpenStudy (anonymous):

a circle with a 60 degreee angle and an arc length of 4 ft, find circumference

OpenStudy (confluxepic):

Welcome to Open Study.

TheSmartOne (thesmartone):

\[\frac{ 60 }{ 360 }\times 2\times \pi \times 4\]

TheSmartOne (thesmartone):

\[\frac{ 1 }{ 6 }\times 8\times \pi\]

TheSmartOne (thesmartone):

\[\frac{ 4 }{ 3 }\pi\]

OpenStudy (confluxepic):

The Smart One is correct.

TheSmartOne (thesmartone):

the circumference is \[\frac{ 4 }{ 3 }\pi\]

OpenStudy (anonymous):